OpenStudy (anonymous):

Find the inverse function of f. f(x) = 10 + √7- 4x

OpenStudy (anonymous):

y=10+squrt7-4x ok to find the inverse you just flip the x and y and solve for Y x=10+squrt7-4y x-10=squrt7-4y x-10-squrt7=4y (x-10-squrt7)/4=y